# Strong Lensing Effect and Quasinormal Modes of Oscillations of Black Holes in $\boldsymbol{f(R,T)}$ Gravity Theory

### v1

@@ III.2 Observables
Figure 7: The outermost relativistic Einstein rings for SgrA* by considering it as a minimally coupled $f(R,T)$ BH, for different values of model parameters $\beta$. The left plot is for $\omega=0$, the middle plot is for $\omega=1/3$ and the right one is for $\omega=0$. The black dashed rings correspond to Schwarzschild BHs.
It should be noted that $k=1$ corresponds to the outermost Einstein ring. We depict the outermost Einstein ring of the supermassive BH SgrA* located at the center of the Milky Way, by modeling it as a minimally coupled $f(R,T)$ gravity BH for different values of $\beta$. The mass $M$ and the distance $d_{ol}$ from the Earth of SgrA* are taken as $4.3\times 10^{6}M_{\odot}$ and $8.35$ kpc respectively [93, 92, 140, 142, 141]. The rings are portrayed for $\omega=0$, $\omega=1/3$ and $\omega=-2/3$ separately in Fig. 7, for different values of parameter $\beta$. It is observed that relative to each respective field, the radii of rings decrease as $\beta$ increases.
@@
Subsequently, we proceed to estimate these three key lensing observables in the strong-field regime, as their evaluation reveals the nature of lensing BHs. To do this, considering the aforementioned mass and distance of SgrA*, the numerical values of the lensing observables $\vartheta_{\infty}$, $s$ and $r_{mag}$ for the BH are computed and summarized in Tables 4 and 5. To analyze their behaviors in the $f(R,T)$ BHs spacetime, the variations of these observables with respect to the modification parameter $\beta$ are illustrated in Fig. 8. It is evident that all three observables decrease with increasing $\beta$ for all three scenarios, $\omega=0$, $\omega=1/3$ and $\omega=-2/3$. For a given $\beta$, the observables attain smaller values in the presence of a dust field compared to a radiation field around it when $\beta\geq-12.23$. Below this critical limit of $\beta$, the values of all three observables exceed those corresponding to the radiation field. Further, $s$ exhibits a relatively steep decline with increasing $\beta$ for $\omega=0$, followed by a very slower variation, whereas for $\omega=1/3$, the decrease is more gradual before showing nearly constant behavior at higher values of $\beta$. On the other hand, for $\omega=-2/3$, although all three observables initially follow a similar trend as in the cases of $\omega=0$ and $\omega=1/3$ from $\beta=-13.7088$, their behavior deviates beyond $\beta\approx-9.8$. Specifically, the observable $\vartheta_{\infty}$ exhibits a sudden decline, while $s$ and $r_{\text{mag}}$ undergo steep growth upto $\beta=-9.4381$. Notably, good results for these observables are obtained only within this range of $\beta$. Consistent with the horizon structure analysis, though not exactly, the observables approach their corresponding GR values at $\beta=-16.7551$, $\beta=-18.8495$ and $\beta=-13.7088$ for the three respective fields and show deviations from GR as $\beta$ increases.
@@
Table 4: Estimation of numerical values of characteristic strong lensing observables by considering supermassive BH SgrA* as a minimally coupled $f(R,T)$ gravity BH for $\omega=0$ and $\omega=1/3$.
    $\omega=0$ $\omega=1/3$
$\beta$ $\vartheta_{\infty}(\mu\text{as})$ s $(\mu\text{as})$   $r_{\text{mag}}$ $\beta$     $\vartheta_{\infty}(\mu\text{as})$ s $(\mu\text{as})$   $r_{\text{mag}}$
SBH 26.5972 0.033239 6.77155 SBH 26.5972 0.033239 6.77155
-16.7551 26.5972 0.033245 6.77155 -18.8495 26.5972 0.033245 6.77155
-13.7088 23.1238 0.020693 $6$13.7088 22.0691 0.017165 6.24005
-12.23 21.1553 0.015319 6.08105 -12.23 21.1155 0.012752 6.07353
-12 20.2849 0.011467 5.90433 -12 20.6717 0.012676 5.98627
-9.4381 17.9771 0.004636 5.26914 -9.4381 19.3424 0.008827 5.68459
-8 17.2371 0.002949 4.97465 -8 18.8598 0.007126 5.54435
0 15.3559 0.001585 3.90956 0 17.4120 0.003259 5.04722
7 14.7449 0.000717 3.37720 7 16.8619 0.002533 4.80724
Table 5: Estimation of numerical values of characteristic strong lensing observables by considering supermassive BH SgrA* as a minimally coupled $f(R,T)$ gravity BH for $\omega=-2/3$.
$\beta$ $\vartheta_{\infty}(\mu\text{as})$ s $(\mu\text{as})$   $r_{\text{mag}}$ $\beta$     $\vartheta_{\infty}(\mu\text{as})$ s $(\mu\text{as})$   $r_{\text{mag}}$
SBH 26.5972 0.033239 6.77155 -11 15.0242 0.001132 3.6522
-13.7088 26.5972 0.041883 6.77592 -10 13.6885 0.000009 2.12064
-12 17.7773 0.004113 5.19698 -9.7 12.5422 0.000425 3.27039
-11.5 16.1008 0.002314 $4$9.4381 9.42328 0.007979 9.14591
@@
Figure 8: Behaviors of three lensing observables with $\beta.$ The left plot shows the variation of the angular position of the asymptotic relativistic images $\vartheta_{\infty}$, the middle one is for the variation of the angular separation between the outermost and asymptotic relativistic images $s$ and the right plot depicts the variation of relative magnification $r_{mag}$ with $\beta$.
@@ IV Quasinormal Modes of Oscillations
Figure 9: Convergence plots of the different order WKB QNMs of the minimally coupled $f(R,T)$ gravity BHs for $\beta=0.1$. These are applicable for other values of $\beta$ also.
@@
Figure 10: Amplitudes of 3rd order WKB QNMs versus $\beta$ of the minimally coupled $f(R,T)$ BHs for $\omega=0$ (left), $\omega=1/3$ (middle) and $\omega=-2/3$ (right) cases with three values of multipole $l$.
From Fig. 10 it is seen that initially amplitude shows a decreasing trend with the model parameter $\beta$ but it rises as $\beta>-10$ for the case of $\omega=0$. Whereas for the radiation case with $\omega=1/3$, the trend is almost linear, where amplitude increases as the value of parameter $\beta$ increases. A similar situation is observed for the $\omega=-2/3$ case where the amplitude increases with $\beta$ but in a nonlinear fashion. As usual, for higher multipole, we get a larger amplitude in all three cases. Similarly, we plot the damping part of the QNM in Fig. 11. It is seen from the figure that for $\omega=0$, the damping is not stable initially, it decreases a little at the beginning but then starts increasing and again increasing in an irregular manner as $\beta$ is increased. For higher values of $\beta>-8$, it is seen that the damping increases gradually and smoothly which slows down for higher positive values of $\beta$. Further, for higher $l$, the increase in the damping is less compared to lower $l$ with varying degrees depending on the value of $\beta$. For the case of $\omega=1/3$, initially there is a drastic difference in the damping for different $l$, which decreases slightly up to $\beta=-1$. As $\beta$ exceeds $-1$, the damping curves merge very close for all $l$ values and show a smooth increasing pattern with increasing $\beta$. For $\omega=-2/3$, there is a larger variation of damping for $l=1$ compared to the higher multipoles, where the damping initially increases then remains more or less the same with increasing $\beta$ depending on $l$ values.
@@
Figure 11: Dampings of 3rd order WKB QNMs versus $\beta$ of the minimally coupled $f(R,T)$ BHs for $\omega=0$ (left), $\omega=1/3$ (middle) and $\omega=-2/3$ (right) cases with three values of multipole $l$.
In Table 6, we tabulate up to 4th order WKB QNMs for the minimally coupled $f(R,T)$ BHs solution (11) along with the associated errors.
@@
Table 6: Up to 4th order WKB QNMs of the BHs solution (11) for $l=1,$ $2$, $n=0$ and for different model parameters $\omega$ and $\beta$. The approximate errors ($\Delta_{3}$) associated with the WKB QNMs have also been computed.
Multipole $\omega$ $\beta$ 2nd order WKB QNMs 3rd order WKB QNMs 4th order WKB QNMs $\Delta_{3}$
Schwarzschild ($l=1$) - - 0.294510 -0.107671i 0.291114 -0.098001i 0.292954 -0.097386i $2.4290\times 10^{-3}$
$l=1$ 00 $-16.0$ 0.325157 -0.097522i 0.325396 -0.098314i 0.325690 -0.098225i $3.5643\times 10^{-4}$
00 $-5.0$ 0.486328 -0.135807i 0.484093 -0.127572i 0.434974 -0.141978i $2.3687\times 10^{-2}$
00 $1.0$ 0.603433 -0.227997i 0.595620 -0.206432i 0.569958 -0.215726i $1.7825\times 10^{-2}$
00 $10.0$ 0.689967 -0.304221i 0.675111 -0.268832i 0.613410 -0.295873i $3.6511\times 10^{-2}$
Schwarzschild ($l=2$) - - 0.483977 -0.100560i 0.483211 -0.096805i 0.483647 -0.096718i $5.4543\times 10^{-4}$
$l=2$ 00 $-16.0$ 0.502966 -0.096764i 0.503056 -0.097229i 0.503186 -0.097203i $1.4964\times 10^{-5}$
00 $-5.0$ 0.803066 -0.130028i 0.802560 -0.126866i 0.791249 -0.128679i $5.9403\times 10^{-3}$
00 $1.0$ 0.990501 -0.212086i 0.988736 -0.203686i 0.989223 -0.203586i $1.4987\times 10^{-3}$
00 $10.0$ 1.120870 -0.277783i 1.117410 -0.263481i 1.119110 -0.263082i $2.5806\times 10^{-3}$
$l=1$ $1/3$ $-9.0$ 0.316522 -0.097246i 0.324043 -0.119474i 0.360770 -0.107311i $2.2634\times 10^{-2}$
$1/3$ $-5.0$ 0.342608 -0.095205i 0.349209 -0.116750i 0.384425 -0.106055i $2.1598\times 10^{-2}$
$1/3$ $1.0$ 0.384381 -0.097927i 0.382510 -0.090304i 0.343880 -0.100448i $1.9204\times 10^{-2}$
$1/3$ $9.0$ 0.431236 -0.109279i 0.429426 -0.101899i 0.393538 -0.111192i $1.7961\times 10^{-2}$
$l=2$ $1/3$ $-9.0$ 0.522599 -0.091364i 0.524122 -0.099710i 0.532904 -0.098066i $5.6634\times 10^{-3}$
$1/3$ $-5.0$ 0.567216 -0.089682i 0.568563 -0.097839i 0.577116 -0.096389i $5.4240\times 10^{-3}$
$1/3$ $1.0$ 0.637824 -0.092550i 0.637413 -0.089669i 0.628945 -0.090876i $4.5136\times 10^{-3}$
$1/3$ $9.0$ 0.714410 -0.104229i 0.714010 -0.101451i 0.706078 -0.102591i $4.2405\times 10^{-3}$
$l=1$ $-2/3$ $-6.0$ 1.205080 -0.584214i 1.167160 -0.501364i 1.096700 -0.533575i $5.9806\times 10^{-2}$
$-2/3$ $-1.0$ 1.407960 -0.585178i 1.369780 -0.486179i 1.781360 -0.373849i $1.4772\times 10^{-1}$
$-2/3$ $1.0$ 1.450180 -0.571179i 1.413240 -0.469531i 2.022270 -0.328126i $2.4505\times 10^{-1}$
$-2/3$ $6.0$ 1.474860 -0.528203i 1.464650 -0.499012i 2.675790 -0.273145i $5.6155\times 10^{-1}$
$l=2$ $-2/3$ $-6.0$ 1.778890 -0.549720i 1.766960 -0.509770i 1.756690 -0.512751i $1.5949\times 10^{-2}$
$-2/3$ $-1.0$ 2.019120 -0.583751i 2.006430 -0.538219i 2.078650 -0.519519i $2.0389\times 10^{-2}$
$-2/3$ $1.0$ 2.075520 -0.583482i 2.062920 -0.536965i 2.171830 -0.510039i $3.7469\times 10^{-2}$
$-2/3$ $6.0$ 2.167270 -0.572230i 2.156060 -0.528214i 2.358880 -0.482799i $8.3120\times 10^{-2}$
@@
As seen from this Table 6, the QNMs are tabulated for the cases of $\omega=0$, i.e. the dust matter case, $\omega=1/3$, i.e. the radiation case and $\omega=-2/3$, i.e. the quintessence case. We also present the QNMs of the Schwarzschild BH for the multipole $l=1$ and $l=2$ for the purpose of comparison. It is seen that the QNMs of the considered $f(R,T)$ BHs have significant deviations from the Schwarzschild case. As already seen from Fig. 10, a common observation is that with increasing multipole, the amplitude increases. Similar is the trend for the parameter $\beta$ where the amplitude shows an increasing trend. The damping part also shows the same trend with respect to $\beta$ as discussed from Fig. 10 analysis. The associated errors have been computed and come out to be in the range of $10^{-5}$ to $10^{-1}$ depending on the value of $\beta$, $l$ and $\omega$. In general, a specific lower value of $\beta=-16$ and higher $l$ values give lower errors. Whereas for $\omega=-2/3$, the errors are maximum. It needs to be mentioned that the error $\Delta_{3}$ is calculated as [25]

### v2

@@ III.2 Observables
Refer to caption    Refer to captionRefer to caption
Figure 9: The outermost relativistic Einstein rings for Sgr A* by considering it as a minimally coupled $f(R,T)$ BH, for different values of model parameters $\beta$ and $c_{2}$ respectively. The upper row is for changing $\beta$ with $c_{2}=0.5$, and the lower row is for varying $c_{2}$ at $\beta=-9.5$. The left plots are for $\omega=0$, the middle plots are for $\omega=1/3$ and the right two are for $\omega=-2/3$. The black dashed rings correspond to Schwarzschild BHs.
It should be noted that $k=1$ corresponds to the outermost Einstein ring. We depict the outermost Einstein ring of the supermassive BH SgrA* located at the center of the Milky Way, by modeling it as a minimally coupled $f(R,T)$ gravity BH for different values of $\beta$ and $c_{2}$ individually. The mass $M$ and the distance $d_{ol}$ from the Earth of Sgr A* are taken as $4.3\times 10^{6}M_{\odot}$ and $8.35$ kpc respectively [96, 99, 98, 147, 149, 148]. The rings are portrayed for $\omega=0$, $\omega=1/3$ and $\omega=-2/3$ separately in Fig. 9, for different values of parameters $\beta$ and $c_{2}$. It is observed that relative to each respective field, the radii of rings decrease as $\beta$ increases. Similar trend is noticed with increase of $c_{2}$ also.
@@
Subsequently, we proceed to estimate these three key lensing observables in the strong-field regime, as their evaluation reveals the nature of lensing BHs. To do this, considering the aforementioned mass and distance of SgrA*, the numerical values of the lensing observables $\vartheta_{\infty}$, $s$ and $r_{mag}$ for the BH are computed and summarized in Table 4. To analyze their behaviors in the $f(R,T)$ BHs spacetime, the variations of these observables with respect to the modification parameters $\beta$ fixing $c_{2}$ at $0.5$, and with $c_{2}$ taking $\beta=-9.5$ are depicted in Fig. 10. It is evident that all three observables decrease with increasing $\beta$ for all three scenarios, $\omega=0$, $\omega=1/3$ and $\omega=-2/3$. For a given $\beta$, the observables attain smaller values in the presence of a dust field compared to a radiation field around it when $\beta\geq-12.23$. Below this critical limit of $\beta$, the values of all three observables exceed those corresponding to the radiation field. Further, $s$ exhibits a relatively steep decline with increasing $\beta$ for $\omega=0$, followed by a very slower variation, whereas for $\omega=1/3$, the decrease is more gradual before showing nearly constant behavior at higher values of $\beta$. On the other hand, for $\omega=-2/3$, although all three observables initially follow a similar trend as in the cases of $\omega=0$ and $\omega=1/3$ from about $\beta=-13.5$, their behavior deviates beyond $\beta\approx-9.7$. Specifically, the observable $\vartheta_{\infty}$ exhibits a sudden decline, while $s$ and $r_{\text{mag}}$ undergo steep growth beyond $\beta=-9.7$. Notably, good results for these observables are obtained only within this range of $\beta$. The observables approach their corresponding GR values at $\beta=-16.731$, $\beta=-18.8495$ and $\beta=-13.7088$ for the three respective fields and show deviations from GR as $\beta$ increases. Again, $\vartheta_{\infty}$ shows a rapid initial decline with respect to $c_{2}$, followed by a slow variation for all three surrounding fields. However, in the case of $s$ the change related to $\omega=-2/3$ is moderate from the initial point, while for $\omega=0$ and $\omega=1/3$, it exhibits a steep decline before decreasing slowly with $c_{2}$. Further, $r_{mag}$ decreases gradually in an even manner for the dust and radiation fields respectively, whereas it grows slowly as a function of $c_{2}$ in the case of the quintessence field.
@@
Table 4: Estimation of numerical values of characteristic strong lensing observables for different values of parameters $\beta$ and $c_{2}$ by considering supermassive BH Sgr A* as a minimally coupled $f(R,T)$ gravity BH for $\omega=0$, $\omega=1/3$ and $\omega=-2/3$ respectively.
$\beta$ $\vartheta_{\infty}(\mu\text{as})$ s $(\mu\text{as})$   $r_{\text{mag}}$ $c_{2}$     $\vartheta_{\infty}(\mu\text{as})$ s $(\mu\text{as})$   $r_{\text{mag}}$
SBH 26.5972 0.033239 6.77155 0 26.5972 0.033239 6.77592
-16.731 26.5972 0.033245 6.77155 0.2 23.9244 0.012753 6.50475
-13.7088 24.6790 0.025693 6.58678 0.4 21.9233 0.009338 6.22178
$\omega=0$ -12 22.8102 0.017886 6.35276 0.5 21.0943 0.008035 6.07535
-9.4 21.0634 0.009731 6.06947 0.7 19.6843 0.005951 5.77136
0 18.8071 0.009157 5.52895 0.8 19.0777 0.005100 5.61320
7 18.2374 0.008667 5.35067 0.9 18.5238 0.004469 5.45045
SBH 26.5972 0.033239 6.77155 0 26.5972 0.033292 6.77592
-18.8495 26.5972 0.033245 6.77155 0.2 24.5137 0.013801 6.57327
-13.7088 24.0187 0.022997 6.51123 0.4 22.8538 0.010815 6.36418
$\omega=1/3$ -12 23.0824 0.019064 6.39098 0.5 22.1411 0.009619 6.25701
-9.4 22.1228 0.008827 5.68459 0.7 20.8952 0.007637 6.03697
0 20.6021 0.007679 5.97195 0.8 20.3462 0.006804 5.92389
7 20.1400 0.006877 5.87211 0.9 19.8384 0.006052 5.80861
SBH 26.5972 0.033239 6.77155 0 26.5995 0.033810 6.82302
-13.7088 26.5972 0.041883 6.77592 0.2 20.4061 0.020423 6.89216
-12 20.9018 0.010009 6.03828 0.4 16.4323 0.013695 6.96094
$\omega=-2/3$ -11.5 19.47890 0.00828 5.71995 0.5 14.9315 0.011512 6.99518
-11 18.4999 0.009013 5.44296 0.7 12.5612 0.004787 7.06336
-10 17.2128 0.006880 5.04514 0.8 11.6097 0.007357 7.09729
-9.4 16.9101 0.031534 10.3246 0.9 10.7762 0.006459 7.13111
@@
Refer to caption   Refer to caption   Refer to caption
Figure 10: Behaviors of three lensing observables with $\beta$ and $c_{2}$ respectively. The left plots show the variations of the angular position of the asymptotic relativistic images $\vartheta_{\infty}$, the middle plots are for the variations of the angular separation between the outermost and asymptotic relativistic images $s$ and the right plots depict the variations of relative magnification $r_{mag}$. The top two rows show the variations with respect to $\beta$, while the bottom plots show the corresponding variations with $c_{2}$.
III.3 Constraints on the model parameters $\beta$ and $c_{2}$ from EHT observations of Sgr A*
In 2022, the EHT unveiled the groundbreaking observations of the supermassive BH Sgr A* at the center of our galaxy by providing detailed direct measurements such as its emission ring diameter with the surrounding emission features [96]. Such observations serve as powerful means for exploring MTGs by constraining their free parameters, and understanding the strong gravitational field effects [150, 151, 152, 98]. For the Sgr A* BH, the Einstein radius is $\vartheta^{E}=(25.90\pm 1.15)\mu as$ as reported by the EHT collaboration [96, 97]. Using these bounds, we constrain the associated free parameters $\beta$ and $c_{2}$ of our BHs for three considered surrounding fields. For this, we depict $\vartheta^{E}$ as a function of $\beta$ by varying $c_{2}$ over a broad range of $\beta$ as shown in Fig. 11. It is found that the observed EHT constraints on $\vartheta^{E}$ for Sgr A* significantly restrict the parameters of the $f(R,T)$ BHs spacetime as indicated by the blue band in the figure. Accordingly, the EHT observations impose bounds to $\beta$, narrowing its allowable range as $c_{2}$ increases across all three cases. The figure further illustrates that for each value of $c_{2}$, only a limited set of $\beta$ values lie within the blue band.
Refer to caption   Refer to caption   Refer to caption
Figure 11: Constraining of free model parameters $\beta$ and $c_{2}$ of $f(R,T)$ gravity BHs spacetime from the observed EHT Einstein radius $\vartheta^{E}$ of Sgr A*, by plotting $\vartheta^{E}$ as a function $\beta$ with respect to varying values of $c_{2}$ for three respective surrounding fields $\omega=0,1/3$ and 2/3-2/3.
In addition, Fig 12 represents the density plot for $\vartheta^{E}$ with parameters $\beta$ and $c_{2}$ in the x and y axes respectively. We try to impose constraints on these two parameters again from the bounds of $\vartheta^{E}$ of the Sgr A* BH from EHT observations using the density plot. The solid black lines in each plot represent the lower and upper bounds on $\vartheta^{E}$ from the EHT data of Sgr A* BH. Thus the regions between the two lines represent the allowed parameter space whose coordinates ($\beta,c_{2}$) are the allowed values of the model parameters. For smaller values of $c_{2}$, the allowed range of $\beta$ extends towards the positive axis to a good extent as seen from the figure. For lower values of $\beta$, in the region $20<\beta<-10$, the allowed range of $c_{2}$ increases to a good extent in the positive y axis as seen clearly from the figure. However, for further lower values of $\beta<-20$, there exists a narrow allowed parameter region that is not very prominent.
Refer to caption   Refer to caption   Refer to caption
Figure 12: Density plots for constraining the free model parameters $\beta$ and $c_{2}$ of $f(R,T)$ gravity BHs spacetime from the observed EHT Einstein radius $\vartheta^{E}$ of Sgr A* for three respective surrounding fields $\omega=0,1/3$ and 2/3-2/3.
@@ IV Quasinormal Modes of Oscillations
Refer to caption     Refer to caption
Figure 13: Convergence plots of the different order WKB QNMs of the minimally coupled $f(R,T)$ gravity BHs for $\beta=0.1$. These are applicable for other values of $\beta$ also. Here $c_{2}=0.8$ is considered.
@@
Figure 14: Amplitudes of 3rd order WKB QNMs versus $\beta$ of the minimally coupled $f(R,T)$ BHs for $\omega=0$ (left), $\omega=1/3$ (middle) and $\omega=-2/3$ (right) cases with three values of multipole $l$. Here $c_{2}=0.5$ is considered.
From Fig. 14 it is seen that for a fixed value of $c_{2}=0.5$, the amplitude shows an increasing trend with an increase in the $\beta$ parameter for the case of $\omega=0$. Also, for the radiation case with $\omega=1/3$, the trend is almost similar, where amplitude increases as the value of parameter $\beta$ increases. A similar situation is observed for the $\omega=-2/3$ case where the amplitude increases with $\beta$ but in a nonlinear fashion in a lower $\beta$ regime. As usual, for higher multipole, we get a larger amplitude in all three cases.
@@
Figure 15: Dampings of 3rd order WKB QNMs versus $\beta$ of the minimally coupled $f(R,T)$ BHs for $\omega=0$ (left), $\omega=1/3$ (middle) and $\omega=-2/3$ (right) cases with three values of multipole $l$. Here $c_{2}=0.5$ is considered.
Similarly, we plot the damping part of the QNM in Fig. 15 with respect to the parameter $\beta$ for a fixed value of $c_{2}=0.5$. It is seen from the figure that for $\omega=0$, the damping is not stable up to some negative values of $\beta$ depending on the value of $l$. It decreases a little at the beginning but then starts increasing in a nonlinear manner as $\beta$ is increased. For higher values of $\beta>-4$, it is seen that the damping increases gradually and smoothly which slows down for higher positive values of $\beta$. Further, for higher $l$, the increase in the damping is less compared to lower $l$ with varying degrees depending on the value of $\beta$. For the case of $\omega=1/3$, the damping for the case $l=1$ is significantly lower than $l=2$ and $l=3$ case. $l=2$ and $l=3$ cases are close to each other. All three cases show an increase in damping with increasing $\beta$. For $\omega=-2/3$, there is a larger variation of damping for $l=1$ compared to the higher multipoles, where the damping initially increases then remains more or less the same with increasing $\beta$ depending on $l$ values. $l=2$ and $l=3$ plots show greater damping as $\beta$ is increased, compared to $l=1$ case.
Refer to captionRefer to captionRefer to caption
Figure 16: Amplitudes of 3rd order WKB QNMs versus $c_{2}$ of the minimally coupled $f(R,T)$ BHs for $\omega=0$ (left), $\omega=1/3$ (middle) and $\omega=-2/3$ (right) cases with three values of multipole $l$ and $\beta=-2$.
Refer to captionRefer to captionRefer to caption
Figure 17: Dampings of 3rd order WKB QNMs versus $c_{2}$ of the minimally coupled $f(R,T)$ BHs for $\omega=0$ (left), $\omega=1/3$ (middle) and $\omega=-2/3$ (right) cases with three values of multipole $l$ and $\beta=-2$.
As seen from Fig. 16, the amplitude increases with $c_{2}$ for the three multipole cases and three different $\omega$ cases. The damping (Fig. 17), on the other hand, increases for $\omega=0$ and $\omega=-2/3$ cases but decreases for $\omega=1/3$. Here the value of parameter $\beta=-2$ is considered. It is to be noted that for $c_{2}$ values higher than 1.1 and for $\beta=-11.8$, the QNMs are found to be indeterminate. In Table 5, we tabulate up to 4th order WKB QNMs for the minimally coupled $f(R,T)$ BHs solution (11) along with the associated errors.
@@
Table 5: Up to 4th order WKB QNMs of the BHs solution (11) for $l=1,$ 22, $n=0$, $\beta=-2$ and for different model parameters $\omega$ and $c_{2}$. The approximate errors ($\Delta_{3}$) associated with the WKB QNMs have also been computed.
Multipole $\omega$ $\beta$ $c_{2}$ 4th order WKB QNMs 3rd order WKB QNMs 2nd order WKB QNMs $\Delta_{3}$
Schwarzschild ($l=1$) - - - 0.294510 - 0.107671i 0.291114 - 0.098001i 0.292954 - 0.097386i $2.5762\times 10^{-3}$
$l=1$ 0 2.0-2.0 $0.0$ 0.292954 - 0.097386i 0.291114 - 0.098001i 0.294510 - 0.107671i $2.5762\times 10^{-3}$
0 2.0-2.0 $0.1$ 0.306207 - 0.101463i 0.304313 - 0.102094i 0.307813 - 0.112096i $2.8913\times 10^{-3}$
0 5.0-5.0 $0.1$ 0.302563 - 0.099759i 0.300738 - 0.100364i 0.304121 - 0.110088i $2.4242\times 10^{-3}$
0 5.0-5.0 $0.5$ 0.353172 - 0.111042i 0.351468 - 0.111580i 0.354682 - 0.121326i $2.3209\times 10^{-3}$
Schwarzschild ($l=2$) - - - 0.483977 - 0.100560i 0.483211 - 0.096805i 0.483647 - 0.096718i $1.1901\times 10^{-4}$
$l=2$ 0 2.0-2.0 $0.0$ 0.483647 - 0.096718i 0.483211 - 0.096805i 0.483977 - 0.100560i $1.1901\times 10^{-4}$
0 2.0-2.0 $0.1$ 0.505550 - 0.100770i 0.505101 - 0.100859i 0.505891 - 0.104743i $1.9871\times 10^{-3}$
0 5.0-5.0 $0.1$ 0.499527 - 0.099080i 0.499093 - 0.099166i 0.499857 - 0.102943i $5.4445\times 10^{-4}$
0 5.0-5.0 $0.5$ 0.583196 - 0.110337i 0.582782 - 0.110416i 0.583511 - 0.114203i $5.1997\times 10^{-4}$
$l=1$ $1/3$ 2.0-2.0 $0.0$ 0.292954 - 0.097386i 0.291114 - 0.098001i 0.294510 - 0.107671i $2.5762\times 10^{-3}$
$1/3$ 2.0-2.0 $0.1$ 0.297469 - 0.097668i 0.295689 - 0.098256i 0.298964 - 0.107709i $2.6263\times 10^{-3}$
$1/3$ 5.0-5.0 $0.1$ 0.296439 - 0.097333i 0.294646 - 0.0979247i 0.297918 - 0.107368i $2.3329\times 10^{-3}$
$1/3$ 5.0-5.0 $0.5$ 0.312994 - 0.096019i 0.311265 - 0.0965518i 0.314002 - 0.105040i $1.8571\times 10^{-3}$
$l=2$ $1/3$ 2.0-2.0 $0.0$ 0.483647 - 0.096718i 0.483211 - 0.096805i 0.483977 - 0.100560i $1.1901\times 10^{-4}$
$1/3$ 2.0-2.0 $0.1$ 0.491023 - 0.097009i 0.490596 - 0.097094i 0.491337 - 0.100769i $1.8799\times 10^{-3}$
$1/3$ 5.0-5.0 $0.1$ 0.489297 - 0.096684i 0.488866 - 0.096769i 0.489607 - 0.100441i $5.2277\times 10^{-4}$
$1/3$ 5.0-5.0 $0.5$ 0.516199 - 0.095577i 0.515768 - 0.095657i 0.516391 - 0.098960i $4.0751\times 10^{-4}$
$l=1$ 2/3-2/3 2.0-2.0 $0.0$ 0.292954 - 0.097386i 0.291114 - 0.098001i 0.294510 - 0.107671i $2.5762\times 10^{-3}$
2/3-2/3 2.0-2.0 $0.1$ 0.421151 - 0.145533i 0.417478 - 0.146813i 0.423742 - 0.163778i $5.5869\times 10^{-3}$
2/3-2/3 5.0-5.0 $0.1$ 0.398722 - 0.139841i 0.395415 - 0.141010i 0.401333 - 0.156839i $4.1784\times 10^{-3}$
2/3-2/3 5.0-5.0 $0.5$ 0.786104 - 0.300951i 0.774726 - 0.305371i 0.793324 - 0.349878i $1.2654\times 10^{-2}$
$l=2$ 2/3-2/3 2.0-2.0 $0.0$ 0.483647 - 0.096718i 0.483211 - 0.096805i 0.483977 - 0.100560i $1.1901\times 10^{-4}$
2/3-2/3 2.0-2.0 $0.1$ 0.676258 - 0.147199i 0.675357 - 0.147396i 0.676897 - 0.154297i $3.5725\times 10^{-3}$
2/3-2/3 5.0-5.0 $0.1$ 0.644813 - 0.140203i 0.643994 - 0.140381i 0.645424 - 0.146803i $1.0148\times 10^{-3}$
2/3-2/3 5.0-5.0 $0.5$ 1.206000 - 0.309746i 1.202880 - 0.310548i 1.208190 - 0.330497i $3.7180\times 10^{-3}$
@@
As seen from this Table 5, the QNMs are tabulated for the cases of $\omega=0$, i.e. the dust matter case, $\omega=1/3$, i.e. the radiation case and $\omega=-2/3$, i.e. the quintessence case, considering $c_{2}$ and $\beta$ as free parameters. We have also presented the Schwarzschild case for multipole $l=1$ and $l=2$ for the purpose of comparison. It is seen that the QNMs have significant deviations from Schwarzschild case when we introduce a MTG model. As already seen from Fig. 14, a common observation is that with increasing multipole, the amplitude increases. Similar is the trend for the parameter $\beta$ where the amplitude shows an increasing trend. The damping part also shows the same trend with respect to $\beta$ as discussed from Fig. 14 analysis. The associated errors have been computed and come out to be in the range of $10^{-5}$ to $10^{-1}$ depending on the value of $\beta$, $l$ and $\omega$. In general, a specific lower value of $\beta=-16$ and higher $l$ values give lower errors. Whereas for $\omega=-2/3$, the errors are maximum. It needs to be mentioned that the error $\Delta_{3}$ is calculated as [22]